OpenStudy (anonymous):

A father wants to build an octagonal (that is, 8-sided) play pen for his toddler. Each side of the octagon has a length of 3*[sqrt(50)] inches. What is the perimeter of the octagon? (Recall that the perimeter is the sum of each side of a shape or boundary).

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

\[3\sqrt{50} \times 8 = 24\sqrt{50}\]

OpenStudy (acxbox22):

now simplify the radical

OpenStudy (anonymous):

how

OpenStudy (acxbox22):